# Lean-STaR: Learning to Interleave Thinking and Proving

### v3

@@ 4.1 Experimental Setup
We evaluate our method on the MiniF2F benchmark [54]. We use a similar evaluation setting as previous works [50, 44, 51], but use our sampling method instead of best-first search for the evaluation of our thought-augmented theorem proving model as discussed in (§3.3). For each problem, Lean-STaR samples $K=32$ times in parallel with temperature $T=0.7$, and limits the number of times a tactic can be generated to a total of $N=50$. We also limit sampling to $10$ minutes per problem. We choose these settings to resemble the inference budget used in our baselines, which follow previous work [44, 6, 51]. Namely, for best-first search baselines we use beam search to generate the next tactic with $S=32,K=1$ [44, 6, 51]. We do not compare with methods designed for other formal languages such as Jiang et al. [23], Xin et al. [49] since language differences greatly influence the pass rate due to the different tactics and automation. We also do not compare with Lample et al. [25] since they only report $S=32,K=64$ on best-first search, which is approximately equivalent to $S=1,K=512$ for the sampling method, which is too computationally expensive for us.
@@
4.2 Main Results
Our main results are reported in Table 1. Lean-STaR gives a significant improvement over the previous state-of-the-art in Lean. For instance, with a similar inference budget, Lean-STaR achieves 34.8% versus $30.3\%$ in InternLM2 [51] using best-first search and $30.7\%$ in COPRA [39] using GPT-4. With a larger compute budget, Lean-STaR’s performance improves further to 36.1%.
Table 3: Pass rates about InternLM2-Plus-7B on the minif2f-test dataset with Lean. This table shows the pass rates of previous works and our work. $S$ is the number of tactics attempted at each expanded node (assumed to be $1$ in sampling) and $K$ is the total number of search or sampling attempts per problem. In sampling we use temperature 0.7, and in search we use beam search when generating the next tactic. Note that we sample $32$ examples twice when $K=64$ in sampling. In search we follow the setup in [51] and limit each search to $20$ minutes.
Approach Decoding $N$ $K$ $S$ Pass rate
InternLM2-plus-7b [51] (Few-Shot) (from paper) Search 1000 1 32 $43.4\%$
InternLM2-plus-7b [51] (Few-Shot) (reproduced) Search 1000 1 32 $42.6\%$
InternLM2-plus-7b [51] (Few-Shot) Sampling 50 32 1 $40.9\%$
SFT (InternLM2-plus-7b) [51] (Few-Shot) Sampling 50 32 1 $41.3\%$
Lean-CoT (InternLM2-plus-7b) Sampling 50 32 1 $43.4\%$
Lean-STaR (Iter-1) (InternLM2-plus-7b) Sampling 50 32 1 $45.4\%$
InternLM2-plus-7b [51] (Few-Shot) Sampling 50 64 1 $42.2\%$
SFT (InternLM2-plus-7b) [51] (Few-Shot) Sampling 50 64 1 $43.4\%$
Lean-CoT (InternLM2-plus-7b) Sampling 50 64 1 $45.5\%$
Thought augmentation improves theorem proving.
Bootstrapping improves thought-augmented theorem proving.
The second phase of Lean-STaR consists of generating new thoughts and tactics with the current language model, saving those that result in correct proofs, and training on the union of the initial thought-augmented dataset and the saved examples (i.e., expert iteration [34, 52, 36]). Refer to Appendix A.1 for details.
4.3 Performance Analysis by Types and Difficulties
@@
Tasks in minif2f-test are manually formalized from Olympiad type problems, drawn from multiple sources including AIME, AMC, IMO problems, and problems from the MATH dataset [21]. These problems can have different levels of difficulty and types. Table 2 reports the number of problems successfully proved, partitioned by type and difficulty. We see that Lean-CoT improves performance in solving difficult problems on all categories, especially those from mathematics competitions. On top of these improvements, Lean-STaR’s improvements come mainly in Number Theory.
$S\times K=1$ $13.5\%$ $20.9\%$ $21.7\%$
$S\times K=2$ $18.0\%$ ($+4.5\%$) $22.5\%$ ($+1.6\%$) $24.6\%$($+2.9\%$)
$S\times K=4$ $23.3\%$ ($+5.3\%$) $25.0\%$ ($+2.5\%$) $27.5\%$($+2.9\%$)
$S\times K=8$ $27.5\%$ ($+4.2\%$) $27.0\%$ ($+2.0\%$) $30.7\%$ ($+3.2\%$)
$S\times K=16$ $29.9\%$ ($+2.4\%$) $28.3\%$ ($+1.3\%$) $33.6\%$ ($+2.9\%$)
$S\times K=32$ $30.7\%$ ($+0.8\%$) $29.5\%$ ($+1.2\%$) $34.8\%$ ($+1.2\%$)
$S\times K=64$ $30.7\%$ ($+0.0\%$) $30.3\%$ ($+0.8\%$) $36.1\%$ ($+1.3\%$)
4.4 Search and sampling budget
4.5 Experiments with stronger base model and more data
We instantiate Lean-STaR using a stronger language model (InternLM2-Math-plus-7b [51]), which was released after the experiment above. We follow a similar setup to the previous experiment.
In this experiment, we used $140,000$ thoughts annotated by GPT-4o [32] to fine-tune a model (“Lean-CoT”). Then we performed only one iteration of expert iteration and collected about $60,000$ (proof state, thoughts, next-tactic) pairs in data, named “STaR dataset” $D_{1}$. We further fine-tuned the Lean-CoT model on dataset $D_{1}$ to get the Lean-STaR model.
Our new results are reported in Table 3. We can see that Lean-STaR still gives a significant improvement over the baseline. For instance, Lean-STaR achieves $45.4\%$ versus $39.8\%$ in InternLM-plus using sampling with a similar inference budget and $43.4\%$ using best-first search with more inference budget reported in [51]. This results show that both retrospective rationale generation and expert iteration can improve the theorem-proving capabilities on a stronger base model.
4.6 Experiments on expert iteration without CoT
Table 5 shows the result of expert iteration without CoT (i.e., using (state, tactic) pairs only) as well as the result of Lean-CoT and Lean-STaR. Expert iteration alone achieves 43.0%, which is less than Lean-STaR (45.4%). This shows that Lean-STaR’s performance gains do not only come from the use of expert iteration.
Approach # (Continual) Training Data Pass@32
InternLM2-plus-7b (Few-Shot) - $40.9\%$
SFT $231,240$ $41.3\%$ $+0.4\%$
Lean-CoT $140,000$ $43.4\%$ $\mathbf{+2.1\%}$
Lean-STaR $60,000$ $45.5\%$ $\mathbf{+2.1\%}$
Expert iteration (SFT) $60,000$ $43.0\%$ $+1.7\%$
5 Conclusion & Limitations
In this paper, we presented Lean-STaR, a novel approach that significantly enhances the theorem-proving capabilities of language models in formal mathematics by integrating Chain-of-Thought (CoT) rationales into each proof step. Our method begins with generating synthetic rationales using ground-truth tactics retrospectively, followed by fine-tuning the language model to generate these rationales and predict subsequent tactics, resulting in the Lean-CoT model. We further improved this model using expert iteration, fine-tuning it on correct proofs it samples and verifies using the Lean solver. Our contributions include the introduction of the first thought-augmented theorem proving dataset, demonstrating that expert iteration can further improve performance, and achieving new state-of-the-art results on the miniF2F-test benchmark, increasing the pass rate from $30.3\%$ to $36.1\%$. These advancements are not only about improving the accuracy of automated theorem proving, but also offer a scalable and efficient framework for advancing human understanding of mathematics, which may lead to significant impacts in education, scientific discovery, and program verification [11, 24, 37, 4, 16, 31].
Appendix A Additional Experiment Setup
A.1 Lean-STaR Expert Iteration
The second phase of Lean-STaR consists of generating new thoughts and tactics with the current language model, saving those that result in correct proofs, and training on the union of the initial thought-augmented dataset and the saved examples (i.e., expert iteration [34, 52, 36]). We perform two iterations of expert iteration, and provide details on our specific experimental setup below.
The setup of experiment about InternLM2-plus is slightly different. The details are shown in Section  4.5 and Appendix  E.
Appendix B Statistics for our methods as well as the baselines
Approach # (Continual) Training Data Pass@32
InternLM2-Math-7b (Few-Shot) - $28.7\%$
SFT $231,240$ $29.5\%$ $+0.8\%$
Lean-CoT $52,438$ $32.8\%$ $\bm{+3.3\%}$
Lean-STaR (Iter-1) $32,231$ $34.0\%$ $+1.2\%$
Lean-STaR (Iter-2) $19,324$ $\mathbf{34.8\%}$ $+0.8\%$
Appendix C An Example and Explanation of A Formal Proof in Lean
An example of a formal proof in Lean with its visualization is shown in Figure 6, taken from [25]. In the proof, the tactic induction k is is applied to the initial state ($n\leq m\Rightarrow n+k\leq m+k$) and the ITP converts the current state to subgoals case 0 $\land$ case ih: $n\leq m\land n+k\leq m+k\Rightarrow n+(k+1)\leq m+(k+1)$. The case 0: $n\leq m$ is our hypothesis $h_{0}$ so it can be proven by case 0:exact $h_{0}$ tactic. Then, we rewrite the case ih through the nat.succ_le_succ_iff which is a theorem in Lean library means $n\leq m\Leftrightarrow n+1\leq m+1$. After tactics case 0:exact $h_{0}$ and case ih:rw nat.succ_le_succ_iff, the goal state is converted to $n+k\leq m+k$ which is the hypothesis introduced by induction. Therefore, we can complete this proof using tactic exact k_ih.
theorem add_le_add_right (m n k : $\mathbb{N}$) (h0 : n $\leq$ m)
: n + k $\leq$ m + k :=
induction k with
| zero =>
exact h0
| succ k ih =>
rw Nat.succ_le_succ_iff
exact ih
Figure 6: A example proof and its visualization of $n\leq m\Rightarrow n+k\leq m+k$ in Lean, taken from [25]. The induction tactic reduces the initial statement to two subgoals. Then tactics case 0:exact $h_{0}$ and case ih:rw nat.succ_le_succ_iff, case ih:exact k_ih can be applied in turn to complete the proof.
Appendix D Performance Analysis by Types and Difficulties using InternLM2-plus-7b
@@
Appendix E Performance difference of joint training and continue training
@@
Appendix F Retrospective Rationale Generation Prompt of GPT-4
@@
Appendix G Examples of generated Lean proofs with thoughts
@@
theorem aime_1983_p2 (x p : $\mathbb{R}$) (f : $\mathbb{R}$rightarrow $\mathbb{R}$) (h0 : 0 < p $\land$ p < 15) (h1 : p $\leq$ x $\land$ x $\leq$ 15)
(h2 : f x = abs (x - p) + abs (x - 15) + abs (x - p - 15))
: 15 $\leq$ f x :=
cases h1 with h1 h2 -- Reasoning: Given a function \(f\) defined over the real numbers, where \(f(x)\) equals the sum of the absolute values of the differences \(x - p\), \(x - 15\), and \(x - p - 15\), along with conditions on \(p\) and \(x\) that ensure \(x\) lies between \(p\) and \(15\) and \(p\) is positive, we aim to show that \(f(x) $\geq$ 15\). To demonstrate this, we consider all possible cases for the position of \(x\) relative to \(p\) and \(15\), specifically at \(p\) itself, between \(p\) and \(15\), and at \(15\). By evaluating \(f(x)\) for each case, we can establish a lower bound of \(f(x)\) that is greater than or equal to \(15\).
simp [h0, abs_of_nonneg (sub_nonneg.2 h1), abs_of_nonneg (sub_nonneg.2 h2), *] -- Reasoning: The objective is to show that \(f(x)\) is at least 15 for \(x\) within the bounds \(p $\le$ x $\le$ 15\), where \(p\) is positive but less than 15. The provided expression for \(f(x)\) suggests using properties of absolute values and inequalities to simplify and analyze \(f(x)\). By applying simplification rules for absolute values and inequalities, along with the given bounds, we aim to demonstrate the desired inequality \(f(x) $\ge$ 15\).
simp only [abs_of_nonneg] at * -- Reasoning: The objective is to establish an inequality involving an expression with absolute values and a sum, where all terms are functions of \(x\). By recognizing that the absolute value functions are always non-negative, we can simplify the given expression by removing the absolute value symbols, thus focusing on the arithmetic manipulation of the terms to demonstrate the inequality.
simp only [abs_of_nonneg, sub_add_eq_add_sub] at * -- Reasoning: The objective is to demonstrate that the expression ‘15 $\leq$ x - p + |x - 15| + |x - p - 15|‘ holds given the conditions on p‘, x‘, and the function f‘. Since the expression in question involves absolute values which inherently involve non-negativity, manipulating these expressions to highlight their non-negativity or equality is a logical initial step. This involves recognizing that the absolute value of a difference can be rewritten as the difference of absolute values, which simplifies the manipulation of the expression to prove the inequality.
congr 1 -- Reasoning: To establish the inequality ‘15 $\leq$ x + |x - 15| + |x - p - 15| - p‘, we recognize that the left-hand side of the inequality involves a mix of additions and subtractions of real numbers. By introducing a congruence transformation that simplifies operations involving absolute values, we aim to align both sides of the inequality in a form that directly reflects the properties of real numbers under these operations.
simp only [abs_of_nonpos (sub_nonpos_of_le h2), abs_of_nonneg (sub_nonneg_of_le h1)] -- Reasoning: The objective is to prove an inequality involving absolute values and subtractions that aligns the given conditions and constraints. Simplification using the properties of absolute values and the provided bounds can directly lead to the desired conclusion by expressing the absolute values in terms of their underlying expressions and applying the given constraints.
simp [abs_of_pos, *] -- Reasoning: To prove the given inequality involving absolute values, simplification using algebraic properties and given hypotheses will directly lead us to the desired inequality by recognizing the algebraic manipulation and application of the known bounds h1 and h2‘, as well as the non-negativity of p‘.
rw [abs_of_nonpos] -- Reasoning: The objective is to demonstrate that ‘15 $\leq$ 15 + |x - p - 15| - p under the given constraints on p and x‘. Given p < 15‘ and p $\leq$ x $\leq$ 15‘, its apparent that p - 15‘ is non-positive. This observation allows us to apply the property that the absolute value of a non-positive number is equal to its negative value, thereby simplifying the expression on the right-hand side of the inequality.
all_goals linarith -- Reasoning: The goal involves demonstrating a relationship between expressions involving inequalities and subtraction. The simplification of these expressions into a form that directly compares their numerical values can lead to a straightforward application of known inequalities and algebraic properties, demonstrating the inequalitys validity under the given conditions.

### v4

@@ 4.1 Experimental Setup
We evaluate our method on the MiniF2F benchmark (Zheng et al., 2021). We use a similar evaluation setting as previous works (Yang et al., 2023; Welleck & Saha, 2023; Ying et al., 2024), but use our sampling method instead of best-first search for the evaluation of our thought-augmented theorem proving model as discussed in (§3.3). We choose these settings to resemble the inference budget used in our baselines, which follow previous work (Welleck & Saha, 2023; Azerbayev et al., 2023b; Ying et al., 2024). Namely, for best-first search baselines we use beam search to generate the next tactic with $S=32,K=1$ (Welleck & Saha, 2023; Azerbayev et al., 2023b; Ying et al., 2024). We do not compare with methods designed for other formal languages such as Jiang et al. (2022); Xin et al. (2023) since language differences greatly influence the pass rate due to the different tactics and automation. We also do not compare with Lample et al. (2022) since they only report $S=32,K=64$ on best-first search, which is approximately equivalent to $S=1,K=512$ for the sampling method, which is too computationally expensive for us.
4.2 Main Results
Our main results are reported in Table 1. Lean-STaR gives a significant improvement over the base model. For instance, with a similar inference budget, Lean-STaR achieves 34.8% versus $30.3\%$ in InternLM2 (Ying et al., 2024) using best-first search and $30.7\%$ in COPRA (Thakur et al., 2023) using GPT-4. With a larger compute budget, Lean-STaR’s performance improves further to 36.1%.
Table 2: Pass rates about InternLM2-Plus-7B on the minif2f-test dataset with Lean. This table shows the pass rates of previous works and our work. The evaluation setting is the same as Table 1.
Approach Decoding $N$ $K$ $S$ Pass rate
InternLM2-plus-7b (Ying et al., 2024) (from paper) Search 1000 1 32 $43.4\%$
InternLM2-plus-7b (Ying et al., 2024) (reproduced) Search 1000 1 32 $42.6\%$
InternLM2-plus-7b (Ying et al., 2024) Sampling 50 32 1 $40.9\%$
SFT (InternLM2-plus-7b) (Ying et al., 2024) Sampling 50 32 1 $41.3\%$
Lean-CoT (InternLM2-plus-7b) Sampling 50 32 1 $43.4\%$
Lean-STaR (Iter-1) (InternLM2-plus-7b) Sampling 50 32 1 $45.4\%$
InternLM2-plus-7b (Ying et al., 2024) Sampling 50 64 1 $42.2\%$
SFT (InternLM2-plus-7b) (Ying et al., 2024) Sampling 50 64 1 $43.4\%$
Lean-CoT (InternLM2-plus-7b) Sampling 50 64 1 $45.5\%$
Thought augmentation improves theorem proving.
Bootstrapping improves thought-augmented theorem proving.
The second phase of Lean-STaR consists of generating new thoughts and tactics with the current language model, saving those that result in correct proofs, and training on the union of the initial thought-augmented dataset and the saved examples (i.e., expert iteration (Polu et al., 2022; Zelikman et al., 2022; Singh et al., 2023)). Refer to Appendix A.1 for details.
4.3 Experiments with stronger base model and more data
We instantiate Lean-STaR using a stronger language model (InternLM2-Math-plus-7b (Ying et al., 2024)), which was released after the experiment above. We follow a similar setup to the previous experiment.
In this experiment, we used $140,000$ thoughts annotated by GPT-4o (OpenAI, 2023) to fine-tune a model (“Lean-CoT”). Then we performed only one iteration of expert iteration and collected about $60,000$ (proof state, thoughts, next-tactic) pairs in data, named “STaR dataset” $D_{1}$. We further fine-tuned the Lean-CoT model on dataset $D_{1}$ to get the Lean-STaR model.
Our new results are reported in Table 2. We can see that Lean-STaR still gives a significant improvement over the baseline. For instance, Lean-STaR achieves $45.4\%$ versus $40.9\%$ in InternLM-plus using sampling with a similar inference budget and $43.4\%$ using best-first search with more inference budget reported in (Ying et al., 2024). This results show that both retrospective rationale generation and expert iteration can improve the theorem-proving capabilities on a stronger base model.
4.4 Experiments on expert iteration without CoT
Table 3 shows the result of expert iteration without CoT (i.e., using (state, tactic) pairs only) as well as the result of Lean-CoT and Lean-STaR. Expert iteration alone achieves 43.0%, which is less than Lean-STaR (45.4%) in InternLM-plus and achieves 30.7% verus 34.0% in InternLM-base. This shows that Lean-STaR’s performance gains do not only come from the use of expert iteration.
Approach Pass@32 of InternLM-Base Pass@32 of InternLM-Plus
Few-Shot $28.7\%$ $40.9\%$
SFT $29.5\%(+0.8\%)$ $41.3\%(+0.4\%)$
Lean-CoT $32.8\%(\mathbf{+3.3\%})$ $43.4\%(\mathbf{+2.1}\%)$
Lean-STaR $34.0\%(+1.2\%)$ $45.5\%(\mathbf{+2.1}\%)$
Expert iteration (SFT) $30.7\%(+1.2\%)$ $43.0\%(+1.7\%)$
5 Conclusion & Limitations
In this paper, we presented Lean-STaR, a novel approach that significantly enhances the theorem-proving capabilities of language models in formal mathematics by integrating Chain-of-Thought (CoT) rationales into each proof step. Our method begins with generating synthetic rationales using ground-truth tactics retrospectively, followed by fine-tuning the language model to generate these rationales and predict subsequent tactics, resulting in the Lean-CoT model. We further improved this model using expert iteration, fine-tuning it on correct proofs it samples and verifies using the Lean solver. Our contributions include the introduction of the first thought-augmented theorem proving dataset, demonstrating that expert iteration can further improve performance, and achieving new results on the miniF2F-test benchmark, increasing the pass rate from $30.3\%$ to $36.1\%$. These advancements are not only about improving the accuracy of automated theorem proving, but also offer a scalable and efficient framework for advancing human understanding of mathematics, which may lead to significant impacts in education, scientific discovery, and program verification (Carter & Monks, 2013; Kang et al., 2020; Szegedy, 2020; Avigad, 2023; First, 2023; of Sciences, 2023).
Appendix A Additional Experiment Setup
A.1 Lean-STaR Expert Iteration
The second phase of Lean-STaR consists of generating new thoughts and tactics with the current language model, saving those that result in correct proofs, and training on the union of the initial thought-augmented dataset and the saved examples (i.e., expert iteration (Polu et al., 2022; Zelikman et al., 2022; Singh et al., 2023)). We perform two iterations of expert iteration, and provide details on our specific experimental setup below.
The setup of experiment about InternLM2-plus is slightly different. The details are shown in Section  4.3 and Appendix  F.
Appendix B Statistics for our methods as well as the baselines
Approach # (Continual) Training Data Pass@32
InternLM2-Math-7b (Few-Shot) - $28.7\%$
SFT $231,240$ $29.5\%$ $+0.8\%$
Lean-CoT $52,438$ $32.8\%$ $\bm{+3.3\%}$
Lean-STaR (Iter-1) $32,231$ $34.0\%$ $+1.2\%$
Lean-STaR (Iter-2) $19,324$ $\mathbf{34.8\%}$ $+0.8\%$
Appendix C Data Leakage
A risk of using GPT-4 for generating thought annotations is data leakage. Sine miniF2F has been used as a dataset for formal theorem proving for a while, it is possible that miniF2F dataset is already seen by GPT-4 or internLM during pre-training process. However, there are several reasons to believe that data leakage is not likely to be true.
First, our experimental setting (fine-tuning on Mathlib, evaluating on miniF2F) follows a widely used experimental setup in benchmark evaluations in neural theorem proving and InternLM was also evaluated on miniF2F. Therefore, we believe that InternLM has not been exposed to miniF2F.
Also, we observed that most of the proofs generated by Lean-STaR are completely different from the manually written proofs in the miniF2F test dataset. Table 5 shows the analysis of proof generated by Lean-STaR on miniF2F test dataset for Lean. From Table 5 we can see that almost all proofs generated by LeanSTaR are different from the proofs mentioned in the miniF2F. If we set aside straightforward simple cases, there is only one proof generated by LeanSTaR is the same as the proofs mentioned in the miniF2F.
Table 5: Analysis of proof generated by Lean-STaR on miniF2F test dataset for Lean. Similar to the analysis in Thakur et al. (2024).
Proofs found in miniF2F-test Proofs NOT in miniF2F Total
Single-Tactic Simple Proofs Two-Tactic Proofs Longer OR Complex Proofs Total
Tactics Used linarith norm_num nlinarith two tactics > 2 tactics OR 1 tactic multi-args sorry
Proof Count 11 12 2 18 39 82 162 244
Exact Match Successful Proof Count 2 4 0 0 1 7 0 7
1st Tactic Match Successful Proof Count 2 4 0 0 1 7 0 7
Distinct Successful Proof Count 5 7 2 14 19 47 43 90 / 112 (80.36%)
Distinct Successful Proof Count ex Single-Tactic - - - 6 12 18 20 38
All Successful Proof Count 11 12 2 17 22 64 48 112
Appendix D An Example and Explanation of A Formal Proof in Lean
An example of a formal proof in Lean with its visualization is shown in Figure 5, taken from (Lample et al., 2022). In the proof, the tactic induction k is is applied to the initial state ($n\leq m\Rightarrow n+k\leq m+k$) and the ITP converts the current state to subgoals case 0 $\land$ case ih: $n\leq m\land n+k\leq m+k\Rightarrow n+(k+1)\leq m+(k+1)$. The case 0: $n\leq m$ is our hypothesis $h_{0}$ so it can be proven by case 0:exact $h_{0}$ tactic. Then, we rewrite the case ih through the nat.succ_le_succ_iff which is a theorem in Lean library means $n\leq m\Leftrightarrow n+1\leq m+1$. After tactics case 0:exact $h_{0}$ and case ih:rw nat.succ_le_succ_iff, the goal state is converted to $n+k\leq m+k$ which is the hypothesis introduced by induction. Therefore, we can complete this proof using tactic exact k_ih.
theorem add_le_add_right (m n k : $\mathbb{N}$) (h0 : n $\leq$ m)
: n + k $\leq$ m + k :=
induction k with
| zero =>
exact h0
| succ k ih =>
rw Nat.succ_le_succ_iff
exact ih
Figure 5: A example proof and its visualization of $n\leq m\Rightarrow n+k\leq m+k$ in Lean, taken from (Lample et al., 2022). The induction tactic reduces the initial statement to two subgoals. Then tactics case 0:exact $h_{0}$ and case ih:rw nat.succ_le_succ_iff, case ih:exact k_ih can be applied in turn to complete the proof.
@@ Appendix D An Example and Explanation of A Formal Proof in Lean
sectionPerformance Analysis by Types and Difficulties Tasks in minif2f-test are manually formalized from Olympiad type problems, drawn from multiple sources including AIME, AMC, IMO problems, and problems from the MATH dataset (Hendrycks et al., 2021). These problems can have different levels of difficulty and types. Table 6 reports the number of problems successfully proved, partitioned by type and difficulty. We see that Lean-CoT improves performance in solving difficult problems on all categories, especially those from mathematics competitions. On top of these improvements, Lean-STaR’s improvements come mainly in Number Theory.
D.1 Performance Analysis by Types and Difficulties using InternLM2-plus-7b
@@
Appendix E Comparison between search method and sampling method
Table 8: Comparison between search method and sampling method. We use sampling method with hyperparameters $N=50\ \&\ S=1\ \&\ K=32$ and BFS method with $N=50\ \&\ S=32\ \&\ K=1$. All sampling decoding in the paper uses a temperature of 0.7. We use BFS to denotes Best-First Search.
Approach BFS (Sampling) BFS (Beam Search) Sampling
Tactic Prediction in Proving BFS BFS Sampling
Token Decoding in Tactics Sampling Beam-Search Sampling
InternLM2-7b (Few-Shot) $29.1\%$ $30.3\%$ $28.7\%$
SFT $29.9\%$ $30.7\%$ $29.5\%$
Lean-CoT $27.0\%$ $25.4\%$ $32.8\%$
Lean-STaR (Iter-1) $29.1\%$ $26.2\%$ $34.0\%$
Lean-STaR (Iter-2) $29.5\%$ $26.2\%$ $34.8\%$
Appendix F Performance difference of joint training and continue training
@@
$S\times K=1$ $13.5\%$ $20.9\%$ $21.7\%$
$S\times K=2$ $18.0\%$ ($+4.5\%$) $22.5\%$ ($+1.6\%$) $24.6\%$($+2.9\%$)
$S\times K=4$ $23.3\%$ ($+5.3\%$) $25.0\%$ ($+2.5\%$) $27.5\%$($+2.9\%$)
$S\times K=8$ $27.5\%$ ($+4.2\%$) $27.0\%$ ($+2.0\%$) $30.7\%$ ($+3.2\%$)
$S\times K=16$ $29.9\%$ ($+2.4\%$) $28.3\%$ ($+1.3\%$) $33.6\%$ ($+2.9\%$)
$S\times K=32$ $30.7\%$ ($+0.8\%$) $29.5\%$ ($+1.2\%$) $34.8\%$ ($+1.2\%$)
$S\times K=64$ $30.7\%$ ($+0.0\%$) $30.3\%$ ($+0.8\%$) $36.1\%$ ($+1.3\%$)
F.1 Search and sampling budget
### State
a b n : $\mathbb{N}$
$\vdash$ ¬ (n $\neq$ 0 $\land$ a $\neq$ 0) $\leftrightarrow$ n = 0 $\lor$ a = 0
### Tactic
“‘lean4
simp only [not_and_or, ne_eq, not_not]
### Reasoning
Figure 6: Examples of (input, output) pairs of retrospective rationale generation with GPT-4. The full prompt is given in Appendix G. See §3.2.2 for the detailed explanation.
Appendix G Retrospective Rationale Generation Prompt of GPT-4
@@
Appendix H Examples of generated Lean proofs with thoughts